OpenStudy (anonymous):

Find the 30th term in the geometric sequence if t{10}=(512) and t{15}=(16384). HELP!!!

OpenStudy (anonymous):

hi there

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

well it so happens that 512 and 16384 are both powers of 2

OpenStudy (anonymous):

2^9 = 512, and 2^14 = 16384, so the sequence appears to be that t{n} = 2^{n-1}

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

ohhh... kind of.

OpenStudy (anonymous):

so a = 2?

OpenStudy (anonymous):

what formula are you using for the geometric sequence? (i'm not sure what 'a' is in your terminology)

OpenStudy (anonymous):

cos I'm following the tn = a*r^n-1

OpenStudy (anonymous):

so in this case a is just 1

OpenStudy (anonymous):

and r?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

the nth term is 2 to the power of n-1

OpenStudy (anonymous):

so what's the 30th term?

OpenStudy (anonymous):

it's gonna be t30 = 1(2)^30-1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ohhh..

OpenStudy (anonymous):

btw

OpenStudy (anonymous):

do I ever use this: t10 = 512 t15 = 16384 tn = a*r^n-1 t10 = a*r^10-1 512 = a*r^9

OpenStudy (anonymous):

tn = a*r^n-1 t15 = a*r^15-1 16 384 = a*r^14

OpenStudy (anonymous):

yeah, that's the general way to solve this problem. I was just taking a shortcut since I happened to recognize both of those numbers as being powers of 2, and you need a calculator to take the 14th root of something. since I happened to recognize the numbers it was easy to solve without a calculator.

OpenStudy (anonymous):

Divide 16384 ar^14 ------ = ------- 512 ar^9 32 = r^5 +/- 6.4 = r

OpenStudy (anonymous):

ohhh, what if I want to use this method?

OpenStudy (anonymous):

sub 6.4 for r in (1): 512 = ar^9 512 = a(6.4)^9 512 = 180a --- ----- 180 180 2. 84 = a

OpenStudy (anonymous):

well 16384/512 is 32, and they are 5 terms apart. that means something to the 5th power is equal to 32, i.e. r^5 = 32. The easiest way to solve this then is to take the log of both sides: 5 log(r) = log(32) log(r) = log(32)/5 r = e^(log(32)/5) r = 2

OpenStudy (anonymous):

sub -6.4 for 4 in (1) 512 = ar^9 512 = a(-6.4)^9 512 = -180a --- ----- -180 -180 - 2. 84 = a

OpenStudy (anonymous):

Since a = 2.84 & r = 6.4, the first three terms are: 2. 84, 18.18, 116. 33, ... Since a = -2.84 & r = -6.4 the first three terms are: -2.84, 18.18, -116. 33, ...

OpenStudy (anonymous):

we appear to be working on different problems

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

You wrote: 32=r^5 r=6.4 This is wrong, how did you get that value of r?

OpenStudy (anonymous):

ohhh.. how do I do that?

OpenStudy (anonymous):

Oh wait you just divided. That's not right. r to the 5th power is 32, not r times 5.

OpenStudy (anonymous):

Do you know how to use logarithms?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

are you allowed to use a calculator?

OpenStudy (anonymous):

yes, but do you know how to do it by paper-&-pencil?

OpenStudy (anonymous):

I'd like how to know both methods.

OpenStudy (anonymous):

There's no easy way to do it with pencil because you have to take the 5th root of 32, which you need to do with logarithms

OpenStudy (anonymous):

ohh, okay. so by using the calculator?

OpenStudy (anonymous):

take the 5th root of 32

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

hello?